OpenStudy (anonymous):

ammy reads the same number of pages from his book each night. The expression 228 – 12x shows how many pages he has left to read on any given night. What does the number 12 in the second term represent?

OpenStudy (asnaseer):

Think of it this way - lets say you had a book with 300 pages in it and you managed to read 10 pages a night. How many pages would be left to read after one night?

OpenStudy (anonymous):

290

OpenStudy (asnaseer):

correct - and how many left after two nights?

OpenStudy (anonymous):

280

OpenStudy (asnaseer):

correct - now look at how you got these numbers...

OpenStudy (anonymous):

subtracted

OpenStudy (asnaseer):

after one night: 300 - 1x10 = 300 - 10 = 290 after two nights: 300 - 2 x 10 = 300 - 20 = 280

OpenStudy (asnaseer):

what do you think will be the formula to work out how many pages are left to read after "x" nights?

OpenStudy (anonymous):

oh ok thank you

OpenStudy (asnaseer):

yw - I'm glad I was able to help :)